OpenStudy (nadiahossain):

The thickness of a U.S.penny is 5.98*10^-2 inch. What is the height of a stack of 2.5*10^4 pennies? Show Your Work.

OpenStudy (calculusxy):

this is a multiplication question as it requires for you to find the height of \(2.5 \times 10^4\) pennies each with the thickness of \(5.98 \times 10^{-2}\) inch. you can set up your expression as \((2.5 \times 10^4) \times (5.98 \times 10^{-2})\)

OpenStudy (calculusxy):

and then multiply the terms that look similar \((2.5 \times 5.98) \times (10^4 \times 10^{-2})\) i hope you know how to work with scientific notations

OpenStudy (nadiahossain):

I know how to work with scientific notation, but just needed the expression, Thanks!

OpenStudy (bj11684):

I did it

OpenStudy (nadiahossain):

It's 1.495*10^3

OpenStudy (calculusxy):

Correct :)